OpenStudy (anonymous):

(9 – 8)^2 × (24 – 6 ÷ 3)

OpenStudy (saifoo.khan):

Ever heard of "PEDMAS"

OpenStudy (anonymous):

sure, but I'm lazy :(

OpenStudy (saifoo.khan):

Me too. hi 5.

OpenStudy (anonymous):

*Gives Hi-5*

OpenStudy (anonymous):

22

OpenStudy (anonymous):

(9 - 8)^2 * (24 - 6/3) = 1^2 * 24 - 2 = 24 - 2 = 22

OpenStudy (anonymous):

Hope that helped.

OpenStudy (anonymous):

Its such a simple sum, I can't believe you're lazy to solve this.